# Simplifying Triangle Meshes in the Wild

### v1

@@ Abstract.
This paper introduces a fast and robust method for simplifying surface triangle meshes in the wild while maintaining high visual quality. While previous methods achieve excellent results on manifold meshes by using the quadric error metric, they suffer from producing high-quality outputs for user-created meshes, which often contain non-manifold elements and multiple connected components. In this work, we begin by outlining the pitfalls of existing mesh simplification techniques and highlighting the discrepancy in their formulations with existing mesh data. We then propose a method for simplifying these (non-manifold) triangle meshes, while maintaining quality comparable to the existing methods for manifold inputs. Our key idea is to reformulate mesh simplification as a problem of decimating simplicial 2-complexes. This involves a novel construction to turn a triangle soup into a simplicial 2-complex, followed by iteratively collapsing 1-simplices (vertex pairs) with our modified quadric error metric tailored for topology changes. Besides, we also tackle textured mesh simplification. Instead of following existing strategies to preserve mesh UVs, we propose a novel perspective that only focuses on preserving texture colors defined on the surface, regardless of the layout in the texture UV space. This leads to a more robust method for textured mesh simplification that is free from the texture bleeding artifact. Our mesh simplification enables level-of-detail algorithms to operate on arbitrary triangle meshes in the wild. We demonstrate improvements over prior techniques through extensive qualitative and quantitative evaluations, along with user studies.
journal: TOG
@@ 1. Introduction
Mesh simplification creates different levels of detail (LOD) of 3D objects while maintaining their visual fidelity. It plays a critical role in interactive graphics to achieve the required performance on rendering and simulation across various hardware. Its importance has been motivating decades of development with a plethora of solutions, such as the widely used QEM by Garland and Heckbert [1997]. Though these solutions can produce high-quality results when the input is a manifold surface, we demonstrate that they fail to generate acceptable results on a majority of meshes obtained from online repositories (see Figure 1, Figure 2).
Figure 2. Off-the-shelf mesh simplification techniques (e.g., [Garland and Heckbert, 1998]) produce excellent results when the input mesh meets the common assumption of being a single manifold mesh (top green). However, when the input mesh does not meet the input assumption, existing methods often fail to achieve satisfactory results (bottom green). As these “defective” inputs form the majority of human-created meshes, this motivates our work: a mesh simplification method that can handle both cases (blue). Note that for the implementation of [Garland and Heckbert, 1998], we use the version from [Cignoni et al., 2008] with topology preservation.
@@
This issue is caused by the discrepancy between existing problem formulations and real-world situations. Existing LOD techniques often assume the input mesh is manifold [Dey et al., 1999]. This was a realistic assumption when, e.g. in the 90s, the goal was to simplify meshes extracted from isosurfacing methods, such as marching cubes [Lorensen and Cline, 1987]. However, recent advancements in 3D modeling have populated online repositories with a rich amount of manually modeled shapes. In contrast to meshes reconstructed from isosurfacing, a majority of them are non-manifold (e.g., 98.9% on ShapeNet [Chang et al., 2015] are non-manifold, see Figure 3). The significant amount of non-manifold meshes demonstrates the importance of revisiting mesh simplification and its problem formulation based on today’s scenario.
@@
Figure 3. Many meshes one can encounter in online repositories are non-manifolds. For instance, Thingi10k [Zhou and Jacobson, 2016] (a relatively clean data with only 3D-printable shapes) possesses 22.5% non-manifold shapes, ModelNets [Wu et al., 2015] (a dataset of CAD models) contains 54.3% non-manifold objects, and more severely 98.9% of the meshes in the ShapeNet dataset [Chang et al., 2015] (a widely used dataset for machine learning) are non-manifolds.
@@
A desired mesh simplification method must be performant and robust to all types of geometric artifacts. In real-time applications such as virtual reality and video games, these (defective) meshes appear frequently during game time from user interactions (i.e., engineers collaboratively model 3D buildings). Being able to efficiently and robustly simplify all the newly created 3D content on the fly is critical to achieving target performance. Such high performance requirements also exclude expensive alternatives, such as repairing defective meshes into manifolds and performing simplification afterward, which may take minutes to hours.
@@
In this work, we propose a robust method to efficiently simplify triangle meshes that may contain multiple components, non-manifolds, and even the extreme case like triangle soups (Figure 4). Unlike previous methods that assume manifold inputs, we elevate the restriction and allow our approach to work with a soup of triangles. This enables us to handle all triangle meshes one can encounter today. To handle textured meshes, instead of preserving UVs like prior arts, we propose to preserve textures during simplification via successive mapping.
This paper starts with a detailed analysis of the pitfalls of existing mesh simplification strategies and a combination of them (section 4). We then detail our robust mesh simplification method which consists of (1) a construction strategy of simplicial 2-complex tailored for mesh simplification (section 5), (2) a modified quadric error metric with a small change in the metric accumulation strategy that leads to significant improvements in the output quality (section 6), and (3) a novel texture preservation simplification strategy based on successive mapping (section 7). We demonstrate that our method can robustly simplify triangle meshes in the wild and achieve better results than prior arts on qualitative, quantitative, and user study evaluations (section 9). Our work contributes to the emerging toolkit for robust geometry processing, such as [Jacobson et al., 2013; Sawhney and Crane, 2020], by offering a robust mesh LOD technique.
@@ 2. Related Work
Our method belongs to the family of local decimation schemes defined on triangle meshes, aiming to preserve their geometric appearance. We thus focus our discussion on prior techniques closest to ours, and only briefly touch other formulations, such as global optimization and computation-preserving simplification. We omit the discussion on simplifying other shape representations, such as quadrilateral meshes, implicit functions, and point clouds, and refer readers to Luebke [2003] for a more comprehensive discussion.
@@
Mesh simplification has been extensively studied in computer graphics to reduce the resolution of 3D objects while preserving their appearance. Early attempts try to fit low resolution polygons [Jr. and Zyda, 1991], globally remesh the input to a lower resolution [Hoppe et al., 1992; Turk, 1992], or locally remove each mesh element [Schroeder et al., 1992]. Since then, different formulations have been proposed, such as clustering of triangles [Cohen-Steiner et al., 2004; Xu et al., 2024] or vertices [Rossignac and Borrel, 1993; Low and Tan, 1997; Lindstrom, 2000], greedily removing triangles [Hamann, 1994; Gieng et al., 1998] or vertices [Klein, 1998], simplifying a manifold envelope wrapping around the input [Mehra et al., 2009; Gao et al., 2022; Chen et al., 2023b], or optimization with a differentiable renderer [Hasselgren et al., 2021; Deliot et al., 2024; Knodt et al., 2024]. Some of these techniques have also been revisited with learning-based techniques [Potamias et al., 2022; Chen et al., 2023a]. Among them, one of the most popular strategies is the iterative edge collapse [Hoppe, 1996] with the quadric error metric [Garland and Heckbert, 1997]. The popularity of edge collapses is partly due to the high-quality output from the quadric error metric and its efficiency, exhibiting constant time complexity for each edge collapse.
@@
The success of quadric error [Garland and Heckbert, 1997] has been motivating the development of a variety of quadrics for measuring different geometric quantities. Lindstrom and Turk [1998] propose volume and area quadrics to measure volume and area changes, respectively. Garland and Heckbert [1998] introduce the boundary quadric to maintain shape boundaries. Trettner and Kobbelt [2020] consider a mesh element as a sample from a probability distribution and derive the probabilistic quadric to further improve its numerical behavior. Other extensions generalize the quadric error to incorporate surface attributes (e.g., textures) by defining quadric error in a higher dimension [Garland and Heckbert, 1998; Garland and Zhou, 2005] or by adding extra quadric terms derived from a linear approximation of the attribute [Hoppe, 1999]. The optimal choice of metric varies depending on the input geometry and applications. The suite of quadric error metrics provides us with several options to mix-and-match them for different tasks, including the ones beyond mesh simplification such as quadric-inspired surface reconstruction [Kobbelt et al., 2001; Ju et al., 2002; Schaefer and Warren, 2005] and filtering [Vieira et al., 2010; Legrand et al., 2019].
@@
Many of the methods mentioned above are designed for simplifying a single manifold triangle mesh. Another class of mesh simplification generalizes the local decimation method to triangle meshes in the wild, which may contain multiple components or non-manifolds. Garland and Heckbert [1997] reformulate the edge collapse operation as an operation to collapse any vertex pair, meaning one can merge two vertices even if they are not connected by an actual edge. These vertex pairs are computed based on vertex-to-vertex distance (see Figure 12) and optimized with the quadric error metric. Instead of vertex pair collapses, Schroeder [1997] identifies different cases for manifold/non-manifold vertices and removes each vertex based on a distance-to-plane metric. If camera information is given, El-Sana and Varshney [1999] demonstrate that one can incorporate camera information into the simplification, In order to handle any triangle meshes, Popovic and Hoppe [1997] reformulate it as a problem of simplifying simplicial complex. Our method is another instance of topology varying simplification for simplicial complex with high quality geometric approximation and is capable of transferring textures.
Another possible direction is to repair the mesh, i.e. convert it into a single manifold, and then simplify it. However, this depends on a repairing method to produce high quality outputs, which is not always the case (see Figure 5).
Figure 5. Repairing a triangle mesh with [Hu et al., 2020] (third) and then simplifying the repaired model (fourth) depends on a having high quality repaired model. However, this is not always the case. Our method simplifies the input mesh directly and leads to a better result (first).
Alternatively, one can robustly create a manifold “cage” to wrap the triangle mesh and then simplify the manifold wrapper [Nooruddin and Turk, 2003; Chen et al., 2023b]. However, these methods, e.g. Chen et al. [2023b], often suffer from deleting interior components which may be important for interacting with the 3D object (see Figure 6).
Figure 6. Compared to methods that construct a manifold mesh wrapper and then simplify it [Chen et al., 2023b] (middle), our method directly simplifies the input and preserves interior structures (right).
We thus follow the spirit of early attempts to simplify triangle meshes with vertex pair collapses.

### v2

# Simplifying Textured Triangle Meshes in the Wild

@@ Abstract.
This paper introduces a method for simplifying textured surface triangle meshes in the wild while maintaining high visual quality. While previous methods achieve excellent results on manifold meshes by using the quadric error metric, they struggle to produce high-quality outputs for meshes in the wild, which typically contain non-manifold elements and multiple connected components. In this work, we propose a method for simplifying these “wild” textured triangle meshes. We formulate mesh simplification as a problem of decimating simplicial 2-complexes to handle multiple non-manifold mesh components collectively. Building on the success of quadric error simplification, we iteratively collapse 1-simplices (vertex pairs). Our approach employs a modified quadric error that converges to the original quadric error metric for watertight manifold meshes, while significantly improving the results on wild meshes. For textures, instead of following existing strategies to preserve UVs, we adopt a novel perspective which focuses on computing mesh correspondences throughout the decimation, independent of the UV layout. This combination yields a textured mesh simplification system that is capable of handling arbitrary triangle meshes, achieving to high-quality results on wild inputs without sacrificing the excellent performance on clean inputs. Our method guarantees to avoid common problems in textured mesh simplification, including the prevalent problem of texture bleeding. We extensively evaluate our method on multiple datasets, showing improvements over prior techniques through qualitative, quantitative, and user study evaluations.
geometry processing, mesh simplification
article: 181journal: TOGcopyright: acmlicensedjournal: TOGjournalyear: 2025journalvolume: 44journalnumber: 6publicationmonth: 12doi: 10.1145/3763277ccs: Computing methodologies Mesh geometry modelscopyright: acmlicensedjournal: TOGjournalyear: 2025journalvolume: 44journalnumber: 6publicationmonth: 12doi: 10.1145/3763277
@@ 1. Introduction
Mesh simplification creates different levels of detail (LOD) for 3D objects while maintaining their visual fidelity.
Figure 2. Comparison between our textured mesh simplification method and a representative prior technique [Garland and Heckbert, 1998] (using the implementation from [Cignoni et al., 2008]). On single component and manifold inputs, both methods produce excellent results (compare baseline, top center, vs. ours, top right). However, for challenging “wild” inputs, often characterized by non-manifold and multiple components geometry, the baseline approach frequently yields unsatisfactory results (bottom center). In contrast, our method preserves visual fidelity and geometric structure more effectively on these wild meshes (bottom right).
@@
It plays a critical role in interactive graphics to achieve target performance in rendering and simulation across various hardware platforms. Its importance has been motivating decades of development with a plethora of solutions, such as the widely used QEM by Garland and Heckbert [1997]. While these solutions produce high-quality results for manifold surface inputs, we demonstrate that they often fail to generate acceptable results on meshes obtained from online repositories (see Figure 1). This issue is caused by the discrepancy between existing problem formulations and real-world data characteristics. Existing LOD techniques often assume the input mesh is manifold [Dey et al., 1999]. This was a realistic assumption when, for instance during the 1990s, the goal was to simplify meshes generated by isosurfacing methods, such as marching cubes [Lorensen and Cline, 1987]. However, recent advancements in 3D modeling have populated online repositories with a rich amount of manually modeled shapes. In contrast to meshes reconstructed from isosurfacing, the majority of them are non-manifold (e.g., 98.9% on ShapeNet [Chang et al., 2015] are non-manifold, see Figure 3).
@@
Figure 3. Many meshes available in online repositories are non-manifold. For instance, Thingi10k [Zhou and Jacobson, 2016] (a dataset consisting of 3D-printable shapes) includes 22.5% non-manifold shapes, ModelNet [Wu et al., 2015] (a dataset of CAD models) contains 54.3% non-manifold objects, and more severely 98.9% of the meshes in the ShapeNet dataset [Chang et al., 2015] (a widely used dataset for machine learning) are non-manifold.
This prevalence of non-manifold meshes underscores the need to revisit mesh simplification and its problem formulation to better suit contemporary datasets.
@@
A desired mesh simplification method must be performant and robust to all types of geometric artifacts, and can preserve surface attributes (e.g., textures). In real-time applications such as virtual reality and video games, these (defective) meshes appear frequently during game time from user interactions (i.e., engineers collaboratively model 3D buildings). Being able to efficiently and robustly simplify all the newly created 3D content on the fly while preserving attributes is critical to achieving target performance. Such requirements also exclude expensive alternatives that do not preserve attributes (see Figure 6), such as repairing defective meshes into manifolds and performing simplification afterward, which may take minutes to hours.
@@
In this work, we propose a method to simplify any textured triangle meshes that may contain multiple components, non-manifold geometry, and even the extreme case like triangle soups (Figure 4). We elevate the manifold restriction of prior arts to work with a soup of triangles, while guaranteeing to converge to the classic quadric error simplification results on manifold inputs. To handle textured meshes, we successively compute correspondences between the input mesh and the simplified mesh, unlike prior arts that aim at preserving UV coordinates. Our strategy reduces texture distortion and prevents texture bleeding occurs during the simplification process. This combination enables us to handle arbitrary textured meshes one can encounter today, without degraded performance on clean inputs (see Figure 2). We provide a background (Appendix A) and analyze the limitations of existing strategies (section 3). This analysis highlights issues with current tools and motivates revisiting the simplification problem for mesh data in-the-wild.
@@ 2. Related Work
Our method belongs to the family of local decimation schemes defined on triangle meshes, aiming to preserve their surface attributes. We thus focus our discussion on prior techniques closest to ours, and refer readers to Luebke [2003] for a more comprehensive discussion.
@@
Mesh simplification has been extensively studied in computer graphics to reduce the resolution while preserving the appearance of 3D objects. Early attempts tried to fit low resolution polygons [Jr. and Zyda, 1991], globally remesh the input to a lower resolution [Hoppe et al., 1992; Turk, 1992], or locally remove each mesh element [Schroeder et al., 1992]. Since then, different formulations have been proposed, such as clustering of triangles [Cohen-Steiner et al., 2004; Xu et al., 2024] or vertices [Rossignac and Borrel, 1993; Low and Tan, 1997; Lindstrom, 2000], greedily removing triangles [Hamann, 1994; Gieng et al., 1998] or vertices [Klein, 1998], simplifying a manifold envelope wrapping around the input [Mehra et al., 2009; Gao et al., 2022; Chen et al., 2023b], or optimization with a differentiable renderer [Hasselgren et al., 2021; Deliot et al., 2024; Knodt et al., 2024]. Some of these techniques have also been revisited in the context of machine learning-driven simplification [Potamias et al., 2022; Chen et al., 2023a]. Among them, perhaps the most popular strategy is the greedy edge collapse [Hoppe, 1996] with the quadric error metric [Garland and Heckbert, 1997]. The popularity of edge collapses is partly due to the efficiency as each edge collapse exhibits constant time complexity.
Figure 5. Repairing a triangle mesh with [Hu et al., 2020] (third) and then simplifying the repaired model (fourth) depends on having a high quality repaired model, which is not always the case. Our method simplifies the input mesh directly and leads to a better result (first).
Figure 6. Running mesh repairing [Hu et al., 2020] (second column) and then simplifying the mesh with [Garland and Heckbert, 1997] (third column) leads to losses on surface attributes, such as textures, and may suffer from suboptimal results due to inperfect repairing. In contrast, our method directly simplifies the input, leading to better results (fourth column).
Figure 7. Comparing to methods that construct a manifold mesh wrapper and simplify it [Chen et al., 2023b] (middle), our method directly simplifies the input and preserves interior structures (right).
@@
The success of quadric error [Garland and Heckbert, 1997] has motivated a variety of quadrics for measuring different geometric quantities. Lindstrom and Turk [1998] propose volume and area quadrics to measure the squared volume and area changes, respectively. Garland and Heckbert [1998] introduce the boundary quadric to maintain the shape boundary. Trettner and Kobbelt [2020] consider that a mesh element is a sample from a probability distribution and derive the probabilistic quadric to further improve its numerical behavior. Other extensions generalize the quadric error to incorporate surface attributes (e.g., textures) by defining quadric error in a higher dimension [Garland and Heckbert, 1998; Garland and Zhou, 2005] or by linear approximation of the attribute [Hoppe, 1999]. The optimal choice of metric varies depending on the input geometry and applications. The suite of quadric error metrics provides us with several options to mix-and-match them for different tasks, including the ones beyond mesh simplification such as quadric-inspired surface reconstruction [Kobbelt et al., 2001; Ju et al., 2002; Schaefer and Warren, 2005] and filtering [Vieira et al., 2010; Legrand et al., 2019].
@@
Many of the methods mentioned above are designed for simplifying a single manifold triangle mesh. Another class of mesh simplification generalizes the local decimation method to triangle meshes in the wild, which may contain multiple components or non-manifolds. Garland and Heckbert [1997] reformulate the edge collapse operation as an operation to collapse any vertex pair, meaning one can merge two vertices even if they are not connected by an actual edge. These vertex pairs are computed based on vertex-to-vertex distance (see Figure 15) and optimized with the quadric error metric. Schroeder [1997] identifies different cases for manifold/non-manifold vertices and removes each vertex based on a distance-to-plane metric. Popovic and Hoppe [1997] strive for maximizing the compression rate and suggest not optimizing the vertex location. In order to handle any triangle meshes, Popovic and Hoppe [1997] reformulate it as a problem of simplifying simplicial complex. Our method is another instance of topology varying simplification for simplicial complex with high-quality geometric approximation and is capable of transferring textures.
Another possible direction is to repair the mesh, i.e. convert it into a single manifold, and then simplify it. However, this depends on a repairing method to produce high-quality outputs, which is not always the case (see Figure 5) and can lose important surface attributes (see Figure 6). Alternatively, one can robustly create a manifold “cage” to wrap the triangle mesh and then simplify the manifold wrapper [Nooruddin and Turk, 2003; Chen et al., 2023b; Portaneri et al., 2022]. However, these methods, e.g. Chen et al. [2023b], often suffer from deleting interior components which may be important for interacting with the 3D object (see Figure 7). We thus follow the spirit of early attempts to simplify triangle meshes with vertex pair collapses.